OpenStudy (anonymous):

Two positive integers have a product of 32. one integer is twice the other ...what are the integers?

OpenStudy (anonymous):

Let the first integer be x and then the other will be: 2x x*(2x) = 32 \[2x^2 = 32\] Divide by 2, \[x^2 = 16\] Take the root both the sides and find x from it..